OpenStudy (anonymous):

complete the square to write the quadratic equation c(x)= x^2-3/4x+1 in vertex form

OpenStudy (anonymous):

is it \[\large c(x)=x^2-\frac{3}{4}x+1\]?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

lets just find the vertex instead, how about that?

OpenStudy (anonymous):

the first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case is \[-\frac{-\frac{3}{4}}{2}=\frac{3}{8}\]

OpenStudy (anonymous):

so it is going to be \[c(x)=(x-\frac{3}{8})^2+k\] to find \(k\) replace \(x\) by \(\frac{3}{8}\) in the original expression, in other words compute \[c(\frac{3}{8})\]

OpenStudy (anonymous):

would the coordinates of the vertex be (.375, 1.421875)

OpenStudy (anonymous):

lets skip decimals and use fractions instead

OpenStudy (anonymous):

so (3/8, 91/64) ?

OpenStudy (anonymous):

hmmm i think the 91 is wrong

OpenStudy (anonymous):

\[\left(\frac{3}{8}\right)^2-\frac{3}{4}\times \frac{3}{8}+1\] is the first step

OpenStudy (anonymous):

you should get \[\frac{55}{64}\] i believe

OpenStudy (anonymous):

my calculator is still spitting out 91/64 ok so (3/8)^2 = .140625 -3/4(3/8)= -.28125 .140625 + 1 -.28125= 1.421875 or 91/64

OpenStudy (anonymous):

did I do anything incorrectly

OpenStudy (anonymous):

\[\frac{9}{64}-\frac{9}{32}+1\\ \frac{9}{64}-\frac{18}{64}+1\\ \frac{-9}{64}+1\\ \frac{55}{64}\]

OpenStudy (anonymous):

Oh okay, and then after we find the vertex what do we do?

OpenStudy (anonymous):

then write it in vertex form \[c(x)=\left(x-\frac{3}{8}\right)^2+\frac{55}{64}\]

OpenStudy (anonymous):

Thanks!! I understand this now!!